OpenStudy (anonymous):

Solve for x. Round your answer to two decimal places. Show your work for full credit.

OpenStudy (anonymous):

This is as far as i have gotten and idk what to do next. Sin 39= 9/x Sin 39(x)=9 X=9/sin 39